OpenStudy (anonymous):

Venus is about 10^8 kilometers from the Sun. Saturn is about 10^9 kilometers from the Sun. About how many times farther from the Sun is Saturn than Venus? 10^72 10^17 10 1 @abhisar

OpenStudy (abhisar):

I think you have done a mistake in the question. Is it \(\sf 10^8~and~10^9\) or simply 108 and 109 ?

OpenStudy (anonymous):

OH yes I am sorry it is Venus is 10^8 Saturn is 10^9

OpenStudy (abhisar):

Wait..sorry..brb

OpenStudy (anonymous):

okay

OpenStudy (abhisar):

Yes....so i am back

OpenStudy (anonymous):

okay :)

OpenStudy (abhisar):

owkay let's see \(\sf 10^8\) means 1 followed by eight zeros and similarly \(\sf 10^9\) means 1 followed by nine zeroes.

OpenStudy (abhisar):

To calculate how many times saturn is farther from the sun than venus we need to divide distance of saturn by distance of venus. so can u divide \(\sf \huge \frac{10^9}{10^8}\) ?

OpenStudy (anonymous):

10

OpenStudy (abhisar):

\(\sf \huge \frac{10^x}{10^y}=10^{x-y}\) .......

OpenStudy (abhisar):

Exactly ! It will be \(\sf 10^1\) which is 10

OpenStudy (abhisar):

Very good!

OpenStudy (abhisar):

So saturn is 10 times farther to sun than venus

OpenStudy (anonymous):

Thank you for your help! .. for some reason when I was trying to type in 10 it would not let me press post, so that's what was taking soo long lol:)